OpenStudy (anonymous):

can someone explain how much the volume of a cylinder increases when the radius is increased by ten percent

OpenStudy (anonymous):

let radius be 100 cm and find its volume

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then take radius as 110 and find the volume

OpenStudy (anonymous):

but what about the height amount ?

OpenStudy (anonymous):

u will get the volume increased

OpenStudy (anonymous):

tAKE HEIGHT SAME in both .let it be h

OpenStudy (anonymous):

so your saying you want the height to be 110 aswell ?

OpenStudy (amistre64):

V = pi (r)^2 h let r increase by some p% V' = pi (r(1+p))^2 h what is the increase?

OpenStudy (anonymous):

no height will be same in both cases .i.e. h

OpenStudy (amistre64):

V' - V is the amount of increase if im reading it right

OpenStudy (anonymous):

amistre i dont get why you put (1+p) in the equation??

OpenStudy (amistre64):

because when we increase something by a certain percent: thats the result

OpenStudy (amistre64):

r + r(10%) = r(1+10%)

OpenStudy (anonymous):

so when say 1 + 10% u mean that the radius ias 10% bigger like in 1.10

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

*u say

OpenStudy (amistre64):

as such: V' - V pi r^2 (1.1)^2 h - pi r^2 h pi r^2 h [(1.1)^2 - 1] would be the increase factor

OpenStudy (anonymous):

why did you subtract 1 this time amistre ?

OpenStudy (amistre64):

i just pulled out common factors: let a = pi r^2 h a(1.1)^2 - a = a[1.1^2 - 1]

OpenStudy (anonymous):

oh is that suppose to show me that once you take out the whole (100%) ur left with 10%?

OpenStudy (amistre64):

let me try this process: our original volume is: V = pi r^2 h when we increase r by 10% the new Volume is: V' = pi (r(1.1))^2 h the amount of increase is the difference between V' and V: V' - V

OpenStudy (anonymous):

but i thought that maybe you would be doing the volume + something else because your finding an INCREASE??

OpenStudy (anonymous):

oooh nevermind

OpenStudy (amistre64):

spose we have something that is 2, and we increase it to get 5 the amount of increase, k is just: 2 + k = 5, or solving for the increase k = 5-2

OpenStudy (anonymous):

your saying to do the volume plus 10% minus the original volmue which will give me 10%

OpenStudy (anonymous):

yea i get that

OpenStudy (anonymous):

so basically, the volume of the cylinder will increase by ten percent

OpenStudy (amistre64):

V = pi (r)^2 h V' = pi (r(1.1))^2 h V + k = V' k = V' - V k = pi r^2 h (1.1^2 - 1) k = .21 pi r^2 h the volume increases by 21%

OpenStudy (amistre64):

we can generalize this to any percentage of the radius increase as the volume increases by [(1+p%)^2 - 1] as the radius increases by p%

OpenStudy (anonymous):

im sorry but i dont get how u got .21 . i thought k = the 10% difference but i think im wrong

OpenStudy (amistre64):

hmmm, all i did was subtract the old volume from the new volume. lets compare this to an increase in the radius: r + k = r' k = r' - r k = r(1.1) - r k = r(1.1 - 1) k = .1 r , which tells us the the radius increased by 10% as it should of, agreed?

OpenStudy (anonymous):

ill be right backk i have to go the bathroom

OpenStudy (amistre64):

k

OpenStudy (anonymous):

im back.im looking at ur last example

OpenStudy (anonymous):

yes.i agree that the radius increased by 10%

OpenStudy (amistre64):

the process is the same for volume V + k = V' k = V' - V do you follow how we get to pi r^2 h (1.1^2 - 1) by working the same kind of process?

OpenStudy (anonymous):

no because in that example your showing that your multiplying the volumes

OpenStudy (amistre64):

i havent multiplied volumes, ive subtracted them ... then simplified it, the same process that I demonstrated with finding the increase in the radius

OpenStudy (amistre64):